OpenStudy (anonymous):

what is the equation of a circle that has center (-6, 2) and radius 106?

OpenStudy (anonymous):



OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/general-formula-equation-of-circle.png you have the "h", and the "k" plug them in

OpenStudy (anonymous):

ok and than when you solve it gives you (x+6)^2 + (y-2)^2 = 106 right?

OpenStudy (jdoe0001):

yeap \(\large \begin{array}{llll} (x-&\quad h)^2+(y-&k)^2=&r^2\\ &\quad \uparrow&\uparrow&\uparrow\\ &-6&2&\sqrt{106}\\ \hline\\ (x+&6)^2 + (y-&2)^2 = &106 \end{array}\)

OpenStudy (anonymous):

thanksssss <3

OpenStudy (jdoe0001):

yw